### v3

@@ 0.2.1 Age prediction
For the site-target dependence scenario, from four MRI datasets that contain healthy participants [AOMIC, eNKI, CamCAN, and 1000Brains] 200 images were extracted from each site in disjoint age ranges, forcing a site-target dependence. The same proportion of male and female participants was retained in each age range. The unharmonized method obtained a Mean Average Error (MAE) of 6.20 (Table 2), which is in an expected range according to the literature 36. The predictions using the WDH and TTL schemes showed an improvement in the performance of about 2 years, compared with the unharmonized scheme (Table 2). The harmonization scheme that does not use the test labels (No Target) showed the highest error. As expected, the harmonization process removed the age-related signal in the features and therefore the ML model was unable to learn the feature-target relationship and failed to generate accurate predictions. In this case, the model just predicts the mean population age for all individuals, predicting the individuals to be older in the AOMIC and eNKI datasets and younger in the CamCAN and 1000Brains datasets (Figure 1). PrettYharmonize made better predictions, on average, compared with the unharmonized and No Target methods, improving the MAE, R2, and age bias, and without inducing leakage (Table 2). PrettYharmonize’s performance was similar to the two methods that allowed leakage. The individuals’ predictions can be found in the Supplementary Information Figures Supp 1-5.
For the site-target independence scenario, three datasets [eNKI, CamCAN, and SALD] containing healthy controls were used. The same number of images from each dataset was retained in the 18-80 age range, maintaining an equal proportion of males/females. The unharmonized model obtained an MAE of 6.3144, similar to the performance obtained in the site-target dependence scenario (Table 2). Neither PrettYharmonize nor any of the leakage-prone harmonization schemes (WDH and TTL) showed a performance improvement for any of the sites, compared to the unharmonized model (Figure 1(b)). Moreover, the average performance was similar for all the harmonization schemes including the No Target scheme (Table 2). This result suggests that the EoS signal was also discarded by the ML model, as it was not related to the target. Notably, consistent with our hypothesis, the No Target scheme did not remove important biological information, as this biologically relevant variance was presented across all sites.
@@
(b) Age regression on site-target dependence scenario
@@
Metric Unharmonized PrettYharmonize WDH TTL No Target
Site-target dependence scenario
MAE 6.20 4.12 3.82 4.28 15.93
R2 0.81 0.919 0.925 0.912 -0.007
Age Bias -0.43 -0.26 -0.32 -0.23 -0.998
Site-target independence scenario
MAE 6.314 6.306 6.034 6.153 6.036
R2 0.785 0.769 0.803 0.775 0.790
Age Bias -0.341 -0.423 -0.366 -0.319 -0.361
@@ 0.2.2 Sex classification
Two datasets [eNKI and CamCAN] containing healthy controls were used in this experiment were used for the forced site-target dependence. From the first one, 95% of females were retained, whereas for the second dataset only 5%, forcing a site-target dependence. In this case, the age range of the individuals was completely overlapped. The unharmonized scheme obtained a high performance (AUC = 0.97), which is similar to the performance reported in the literature 37. The harmonization schemes that allow leakage (WDH and TTL) and PrettYharmonize did not show improvement compared to the unharmonized scheme (Table 3). This can be mainly because the features carry a strong signal related to the participant’s sex, yielding high performance, even using unharmonized data. Consistent with the age regression experiment, the harmonization scheme without test labels (No Target), removed sex-related information significantly reducing the model’s classification performance (Table 3).
Using the same generated dataset as in the age regression problem for site-target independence, a sex classification experiment was performed. The unharmonized scheme obtained a slightly lower (AUC=0.918) classification performance compared to the sex classification experiment with site-target dependence (Table 3). The harmonization schemes did not show classification improvement compared with the unharmonized model (Table 3). The No Target scheme does not remove target-related variance while harmonizing the features, explaining the similar performance of this model compared to the other schemes.
@@
Metric Unharmonized PrettYharmonize WDH TTL No Target
Site-target dependence scenario
AUC 0.969 0.968 0.975 0.967 0.703
bACC [%] 92.64 92.18 92.10 92.07 63.08
F1 0.923 0.918 0.917 0.917 0.608
Site-target independence scenario
AUC 0.918 0.921 0.913 0.918 0.919
bACC [%] 84.94 85.06 84.64 85.16 84.85
F1 0.851 0.851 0.847 0.852 0.849
@@ 0.2.3 Dementia and mild cognitive impairment classification
For the site-target dependence scenarios, 100 dementia-MCI patients and 10 controls from one site and 100 controls and 10 dementia-MCI on a second site were selected from the ADNI dataset. The unharmonized method obtained an AUC of 0.81, consistent with other findings in the literature 38. PrettYharmonize and the leakage-prone methods showed a slightly higher classification performance than the unharmonized method. As observed before, No Target removed important biological information, jeopardizing the ML model’s performance (Table 4).
For the site-target independence scenario, the same number of control and patients were selected from both sites. In this scenario, all methods obtained a similar classification performance in all metrics (Table 4). Noteworthy, an important performance drop in all schemes is presented, compared with the site-target dependant experiment.
Table 4: Classification performance metrics for dementia-MCI prediction task in site-target dependent scenarios.
Unharmonized PrettYharmonize WDH TTL No Target
Site-target dependence scenario
AUC 0.8131 0.8429 0.8385 0.8381 0.6384
bACC [%] 73.7273 77.2727 76.6364 76.3636 60.1818
F1 0.7371 0.7715 0.7644 0.7622 0.6054
Site-target independence scenario
AUC 0.7092 0.7089 0.7118 0.7103 0.7096
bACC [%] 65.68 65.31 66.01 65.85 66.23
F1 0.6698 0.6659 0.6755 0.6742 0.6794
@@ 0.2.4 Discharge status prediction of septic patients
From the eICU dataset, 20 sites with more than 50 patients were selected. From these selected sites, the “Alive” patients were removed from 10 sites and the Expired patients were removed from the other 10 sites, forcing a site-target relationship. The unharmonized method obtained an AUC of 0.76, slightly lower than the one obtained in 39. However, this difference is expected, as fewer patients were used in our experiments, compared with the reference study. PrettYharmonize obtained an important AUC performance improvement compared with all the benchmarked schemes. No Target scheme removed almost all the relevant information, obtaining a practically by-chance performance (Table 5).
For the site-target independence scenario, the same 20 sites selected were used, but the same number of Alive and Expired patients were retained in each site. All methods obtained the same classification performance in all metrics. The unharmonized method obtained a slightly lower classification performance (0.72 AUC) compared with the site-target dependence scenario (0.76 AUC) (Table 5). Furthermore, PrettYharmonize and the leakage-prone schemes (WDH and TTL) showed a great drop in classification performance, compared with the site-target dependence scenario. Finally, the No Target method did not remove important information while harmonizing the features, obtaining a similar performance as the rest of the benchmarked methods.
@@
Unharmonized PrettYharmonize WDH TTL No Target
Site-target dependence scenario
AUC 0.7655 0.8588 0.7995 0.7897 0.5723
bACC [%] 64.37 66.25 63.39 63.91 51.66
F1 0.4571 0.4910 0.4408 0.4517 0.0921
Site-target independence scenario
AUC 0.7227 0.7101 0.7029 0.6907 0.7198
bACC [%] 66.88 66.14 65.25 64.75 66.42
F1 0.6250 0.6295 0.6133 0.6091 0.6211
@@ Discussion
Combining data from differing acquisitions is an appealing, and sometimes only, option for building ML models as they typically benefit from greater sample sizes. However, it is important to correctly integrate data harmonization methods, like the widely used ComBat, in ML pipelines 19. In this study, we evaluated several ML pipelines incorporating ComBat-based data harmonization using a wide variety of biomedical data from different domains for both classification and regression tasks. Regarding the use of ComBat, feature distribution before and after harmonization is often analyzed to assess the effectiveness of the harmonization process. While this analysis can confirm that the features have been adjusted to exhibit more similar distributions, this analysis alone does not provide sufficient insight into how these adjustments impact the performance of ML models.
@@
For all the evaluated scenarios, the ML pipelines using unharmonized data showed a performance close to the reported in the literature 36, 38, 39, 37. Importantly, for the site-target dependence scenarios, all the unharmonized models showed a better performance compared with the site-target independent scenarios. This is expected as the ML models can to pick EoS signal, which is related to the target in the site-target dependence scenarios, and use it to fraudulently increase the classification performance.
We observed that ComBat-based harmonization struggles to provide benefits when the site and target variables are independent, even when allowing leakage and the target variance was preserved in ComBat modelling. None of the harmonization-based ML pipelines showed performance improvement over the baseline of pooling the data together with site-target independence. This can be explained because the harmonization may be removing a signal that it is also discarded by the ML models, as it is not related to the target. In this case, the benefit of removing the Effect of Site (EoS), did not improve the signal-to-noise ratio of the signal enough to benefit the performance of the ML models. Despite the wide variety of tasks and data from different domains we tested, ComBat-based harmonization does not seem to provide an advantage. Nevertheless, it is possible that our data and task selection, albeit comprehensive, does not include cases where harmonization can be indeed beneficial.
On the other hand, on the site-target dependence senarios, a performance improvement was observed when when applying ComBat and allowing the target variance to be preserved. Problematically, in conventional ways to integrate ComBat in ML pipelines (Whole Data Harmonization and Test Target Leakage), this leads to data leakage as the test labels are used. Whereas not explicitly preserving the target (No Target), the performance was consistently worse. These observations can be explained as ComBat works on the assumption that the site-specific variance and variance of interest are independent. However, this assumption is violated when the site and target are dependent. Consequently, when the site and target are dependent, using ComBat without explicitly preserving the target variance can remove variance related to the target.
The proposed a new method called PrettYharmonize avoids leakage by design, as it relays in the harmonization using pretended target labels and a Stack model, which combines the predictions made with the different harmonized data. In this way, the method can circumvent the need for target values of the test samples. PrettYharmonize was validated on the MAREoS datasets which were specifically devised for this purpose. The solid results obtained in this dataset indicate that the proposed method effectively harmonizes data without compromising model integrity. The method provides a leakage-free pipeline which in our evaluations showed performs at par with leakage-prone pipelines. These findings suggest that PrettYharmonize holds promise for real-world deployment, particularly in contexts where data leakage is a critical concern. Therefore, we recommend PrettYharmonize for future use cases. Overall, our results suggest that future studies should carefully evaluate their ML pipelines and follow reproducible and open science practices for the benefit of the community.
@@
Alternative approaches such as calculating ComBat parameters using phantoms 18 can harmonize data independent of biological variability. By doing so, the location and scale parameters specific to each MRI setup and parameter setting can be accurately estimated and applied to real data. This approach mitigates the risk of inadvertently removing meaningful biological variation during harmonization. However, such approaches are domain-specific and incur challenges such as measuring additional data.
Finally, although our study focused on widely used ComBat-based methods, it is important to acknowledge the existence of other harmonization techniques, such as deep learning-based methods like style-matching generative models or variational autoencoders 4. These approaches may offer promising alternatives, particularly in complex scenarios where traditional methods may fall short though they tend to be data and compute-intensive.
@@
This study has some limitations that should be considered when interpreting the results. First, our analysis focuses primarily on ComBat-based harmonization methods due to their widespread use; however, we did not extensively explore other emerging techniques such as deep learning, optimal transport, or style-matching generative models, which offer different strengths and weaknesses.
Second, the impact of harmonization on feature selection and model interpretability was not deeply explored, warranting further investigation into how these methods influence model behavior in different contexts.
Third, even though in a controlled fashion we simulated somewhat extreme site-target dependence and independence scenarios, it is safe to assume that any real-case scenario will fall in between. Our aim in this study was to empirically demonstrate the problems that may occur while harmonizing the data without factoring in appropriate considerations. Further research should be conducted to better measure the relationship between the site-target dependence degree and the harmonization impact.
Finally, regarding PrettYharmonize, the proposed method is more computationally expensive than the traditional harmonization schemes because in the “pretending” process the data need to be harmonized several times.
@@
Resource availability
Lead contact
Please send any inquiry to the corresponding author, Nicolás Nieto (n.nieto@fz-juelich.de)
Materials availability
No materials were used in this work.
@@
All used MRI datasets are publicly available possibly upon registration. For the eICU dataset, data is publicly available at https://physionet.org/content/eicu-crd/2.0/ after registration. Registration includes the completion of a training course in research with human individuals at https://about.citiprogram.org/ and signing of a data use agreement mandating responsible handling of the data and adhering to the principle of collaborative research.
The library is publicly available at: https://github.com/juaml/PrettYharmonize. The scripts to replicate the experiments and replicate the processing of the datasets are available at: https://github.com/juaml/harmonize_project
@@
To ensure the validity of PrettYharmonize, we benchmarked it in a classification problem using the datasets specifically designed to evaluate harmonization models 3. This MAREoS dataset consists of eight datasets simulating 18 MRI features (cortical thickness, cortical surface area, or subcortical volumes). Four datasets contain a “True” signal and four only contain an Effect of Site (“EoS”) signal related to a binary target. In that sense, an ML model that learns the “EoS” signal can fraudulently achieve a good classification performance. The signal, both the True or EoS, are called “Simple” and “Interactions”, depending on a linear or non-linear effect, respectively. Within each dataset, approximately 1000 samples, coming from 8 sites, were simulated. The datasets are provided as 10 train and test fold pairs. For the dataset containing only the EoS, the methods should be able to remove this effect and the classification performance should be at the chance level, i.e. balanced accuracy (bACC) of 50%. On the other hand, in the dataset with only the True signal, the harmonization models should not degrade the signal, and the bACC is expected to be a high value (bACC $\approx$ 80%).
@@
To empirically compare different harmonization schemes with and without site-target dependence, age regression, and sex classification were performed using MRI data. These targets were used as they are highly reliable and can be easily obtained. For all T1-weighted MR images, Voxel-Based Morphometry was performed using CAT12.8 40 to obtain modulated gray matter (GM) volume, which was then linearly resampled to 8x8x8 mm3 voxels, resulting in 3747 voxels that were used as features. Five datasets were used: Amsterdam Open MRI Collection (AOMIC-ID1000) 41, The Enhanced Nathan Kline Institute (eNKI) 42, Cambridge Centre for Ageing Neuroscience (CamCAN) 43, 1000Brains 44, and the Southwest University Adult Lifespan Dataset (SALD) 45. These datasets were selected as the data within each dataset was acquired only in one site thus avoiding additional confounding. The demographic information of these datasets is presented in Table 6.
@@
0.2.4.1   Forced site-target dependence.
Four datasets, AOMIC-ID1000, eNKI, CamCAN, and 1000Brains, were randomly subsampled in different age ranges forcing a site-target dependence. The subsample was performed to ensure the same amount of subjects by each sex in each dataset (Table 7).
@@
0.2.4.2   Forced site-target independence.

### v4

@@ 0.2.1 Age prediction
For the site-target dependence scenario, 200 images were extracted from each of four MRI datasets containing healthy participants (AOMIC, eNKI, CamCAN, and 1000Brains) in disjoint age ranges, thereby forcing a site-target dependence. The same proportion of male and female participants was retained in each age range. The Unharmonized method achieved a Mean Absolute Error (MAE) of 6.20 (Table 2), which falls within the expected range according to the literature [31]. Predictions using the WDH and TTL schemes showed an improvement in performance of approximately 2 years compared to the Unharmonized scheme (Table 2).
As expected, No Target removed the age-related signal in the features, preventing the ML model from learning the feature-target relationship and resulting in inaccurate predictions. Specifically, the model predicted the mean population age for all individuals, leading to overestimations in the AOMIC and eNKI datasets and underestimations in the CamCAN and 1000Brains datasets (Figure 1).
PrettYharmonize, on the other hand, achieved better predictions on average compared to both the Unharmonized and No Target methods. It improved the MAE, R², and age bias (Pearson’s correlation between the true age and the difference between the predicted and true age) without inducing data leakage (Table 2). Notably, PrettYharmonize’s performance was comparable to the two leakage-prone methods (WDH and TTL). Detailed individual predictions can be found in Supplementary Information Figures Supp 1-5.
For the site-target independence scenario, three datasets (eNKI, CamCAN, and SALD) containing healthy controls were used. The same number of images from each dataset was retained within the 18-80 age range, maintaining an equal proportion of males and females. The model that used Unharmonized data achieved an MAE of 6.314, similar to its performance in the site-target dependence scenario (Table 2).
In this scenario, the average performance was similar across all harmonization schemes, including the No Target scheme (Table 2). This result suggests that the improvement in the signal-to-noise ratio made by removing of the EoS signal was not sufficient to boost ML performance. Consistent with our hypothesis, the No Target scheme did not remove biologically relevant information, as this variance was shared across all sites, and the ML model showed a comparable performance as the other benchmarked schemes.
@@
(b) Age regression on site-target independence scenario
@@
Harmonization Scheme Site-target dependence Site-target independence
MAE R2 Age Bias MAE R2 Age Bias
Unharmonized 6.20 0.81 -0.43 6.314 0.785 -0.341
PrettYharmonize 4.12 0.919 -0.26 6.306 0.769 -0.423
WDH 3.82 0.925 -0.32 6.034 0.803 -0.366
TTL 4.28 0.912 -0.23 6.153 0.775 -0.319
No Target 15.93 -0.007 -0.998 6.036 0.790 -0.361
@@ 0.2.2 Sex classification
In this experiment, two datasets were utilized: eNKI and CamCAN, both containing healthy controls. To enforce site-target dependence, 95% of the females were retained from the eNKI dataset, while only 5% were retained from the CamCAN dataset. The age ranges of the individuals were completely overlapping across the two datasets. The Unharmonized scheme achieved a high performance (AUC = 0.97), consistent with results reported in the literature [53].
The harmonization schemes that allow data leakage (WDH and TTL) and PrettYharmonize did not show any improvement over the Unharmonized scheme (Table 3). This is likely due to the presence of a strong sex-related signal in the features, which enables high performance (AUC = 0.97) even without harmonization. Consistent with the findings from the age regression experiment, the No Target scheme, removed sex-related information, resulting in a significant drop in classification performance (Table 3).
Using the same dataset generated for the age regression problem under site-target independence scenario, a sex classification experiment was conducted. The Unharmonized scheme achieved a slightly lower performance (AUC = 0.918) compared to the site-target dependence scenario (Table 3). None of the harmonization schemes demonstrated improved classification performance relative to the Unharmonized model (Table 3). Again consistent with the age regression experiment, the No Target scheme did not eliminate target-related variance during harmonization, leading to performance similar to the other schemes.
@@
Harmonization Scheme Site-target dependence Site-target independence
AUC bACC [%] F1 AUC bACC [%] F1
Unharmonized 0.969 92.64 0.923 0.918 84.94 0.851
PrettYharmonize 0.968 92.18 0.918 0.921 85.06 0.851
WDH 0.975 92.10 0.917 0.913 84.64 0.847
TTL 0.967 92.07 0.917 0.918 85.16 0.852
No Target 0.703 63.08 0.608 0.919 84.85 0.849
@@ 0.2.3 Dementia and mild cognitive impairment classification
For the site-target dependence scenario, 100 dementia-MCI patients and 10 controls were selected from one site, while 100 controls and 10 dementia-MCI patients were selected from a second site within the ADNI dataset. The Unharmonized method achieved an AUC of 0.81, consistent with findings reported in the literature [30]. PrettYharmonize, WDH, and TTL showed a slight improvement in classification performance compared to the Unharmonized method (Table 4. As observed in previous site-target dependence experiments, the No Target scheme removed biologically relevant information, significantly impairing the ML model’s performance (Table 4).
For the site-target independence scenario, an equal number of controls and patients were selected in each sites. In this case, all benchmarked schemes achieved similar classification performance across all metrics (Table 4). Notably, a consistent performance drop was observed across all schemes compared to the site-target dependence scenario.
Table 4: Classification performance metrics for dementia-MCI prediction task in site-target dependent and independent scenarios.
Harmonization Scheme Site-target dependence Site-target independence
AUC bACC [%] F1 AUC bACC [%] F1
Unharmonized 0.8131 73.7273 0.7371 0.7092 65.68 0.6698
PrettYharmonize 0.8429 77.2727 0.7715 0.7089 65.31 0.6659
WDH 0.8385 76.6364 0.7644 0.7118 66.01 0.6755
TTL 0.8381 76.3636 0.7622 0.7103 65.85 0.6742
No Target 0.6384 60.1818 0.6054 0.7096 66.23 0.6794
@@ 0.2.4 Discharge status prediction of septic patients
From the eICU dataset, 20 sites with more than 50 patients were selected. To enforce a site-target relationship, “Alive” patients were removed from 10 sites, and “Expired” patients were removed from the other 10 sites. The Unharmonized scheme achieved an AUC of 0.76, slightly lower than values previously reported in the literature [32]. This difference is expected, as fewer patients were used in our experiments compared to the referenced study.
PrettYharmonize demonstrated an improvement in AUC performance compared to all benchmarked schemes, achieving an AUC of 0.86. In contrast, and consistent with the previous site-target dependence scenarios, the No Target scheme removed nearly all relevant information, resulting in performance close to chance (AUC = 0.57) (Table 5).
For the site-target independence scenario, an equal number of “Alive” and “Expired” patients were retained across all 20 sites. All methods achieved similar classification performance across all metrics. The Unharmonized method obtained a slightly lower AUC (0.72) compared to the site-target dependence scenario (0.76) (Table 5).
PrettYharmonize and the leakage-prone schemes (WDH and TTL) showed a drop in classification performance compared to the site-target dependence scenario. The No Target method, however, did not remove biologically relevant information during harmonization, achieving performance similar to the other benchmarked methods (AUC = 0.72).
@@
Harmonization Scheme Site-target dependence Site-target independence
AUC bACC [%] F1 AUC bACC [%] F1
Unharmonized 0.7655 64.37 0.4571 0.7227 66.88 0.6250
PrettYharmonize 0.8588 66.25 0.4910 0.7101 66.14 0.6295
WDH 0.7995 63.39 0.4408 0.7029 65.25 0.6133
TTL 0.7897 63.91 0.4517 0.6907 64.75 0.6091
No Target 0.5723 51.66 0.0921 0.7198 66.42 0.6211
@@ Discussion
In many clinical domains, data collected from a single location is insufficient for training generalizable ML models, as large and diverse datasets are needed to identify meaningful feature-target relationships. Combining data from multiple acquisition sites is thus a common and attractive strategy for developing ML models. However, this introduces unwanted variability due to acquisition idiosyncrasies. Data harmonization methods, such as ComBat, are often used to remove site-related variability with promising results in ML applications [6]. Integrating harmonization methods in ML pipelines brings additional challenges. For instance, care must be taken to avoid data leakage by properly separating training and test sets [55, 28, 50]. Other sources of leakage and the effectiveness and practicality of data harmonization in ML workflows remain under-investigated, particularly in scenarios where the target variable and acquisition site are not independent-a situation in which harmonization methods typically struggle. In this study, we evaluated several harmonization schemes using biomedical datasets from various domains, specifically analyzing their performance in cases of both site-target dependence and independence.
In site-target dependence scenarios, we observed performance improvements when applying ComBat with the target specified as a covariate, which preserves its variance (Table 2,4,5. However, using the target as a covariate inevitably leads to data leakage as the target values of the test set are used in both Whole Data Harmonization (WDH) and Test Target Leakage (TTL) schemes. In the WDH scheme, all available data is used to train the harmonization model, and the whole dataset is transformed before splitting it into train and test sets, leading to a form of leakage known as “preprocessing on training and test sets” [55]. Although in the TTL scheme the test data is not used in the harmonization training process, the ComBat model still requires the target values of the test set to correctly transform the test data, a form of “target leakage” [28]. As expected, using ComBat without the target as a covariate (No Target) avoids data leakage but removes biologically relevant information, as evidenced by the marked and consistent drop in the performance across all evaluated scenarios (Tables 2, 3, 4, 5). These results empirically demonstrate the impact of the site-target dependence, which violates ComBat’s assumption that relevant variance, i.e. target variance, is shared across sites.
Conversely, in site-target independence scenarios, despite testing a wide range of tasks and datasets from different domains, none of the harmonization schemes improved performance over the baseline approach of simply pooling the data. This also includes our proposed PrettYharmonize method which, while not being a harmonization method itself, uses a novel way to leverage harmonization without causing data leakage. This lack of performance improvement may be due to the fact that the harmonization process did not sufficiently enhance the signal-to-noise ratio to allow the models to generate better predictions. Nevertheless, we acknowledge that it is possible that our selection of data and tasks, while comprehensive, may not encompass all scenarios where harmonization could prove beneficial.
@@
The ML pipelines pooling data without any harmonization showed performance consistent with values reported in the literature for all evaluated tasks [31, 30, 32, 53], indicating correct application of ML models. As expected, all models using pooled unharmonized data performed better in site-target dependence scenarios, as the ML models could exploit the EoS signal, which in this case is related to the target, illicitly increasing their classification performance.
@@
An alternative approach to training harmonization models is to use phantoms or traveling subjects [22, 5]. This can allow for accurate estimation of location and scale parameters specific to each acquisition setup and parameter setting, which can then be applied to real data. This approach mitigates the risk of inadvertently removing meaningful biological variation during harmonization. However, it is domain-specific and needs additional data collection which is costly as well as accounting for other challenges such as temporal shifts in equipment.
With respect to clinical data, to our knowledge, this is the first study to explore the potential of applying harmonization to this type of data. It is important to distinguish this form of harmonization, which aims to remove EoS from features, from the more common use of the term in the clinical domain, where “harmonization” typically refers to the standardization of feature names and units across sites [4].
The proposed PrettYharmonize method is designed to avoid data leakage. It first harmonizes the data using pretended labels and uses the harmonized output to learn a prediction model. The main idea behind PrettYharmonize is learning subtle differences caused by use of correct or incorrect labels. This design ensures that the harmonization process remains confined to the training phase and circumvents the need for actual target values from the test samples, ensuring a leakage-free ML pipeline. Several key aspects of PrettYharmonize deserve emphasis. First, it is built around the neuroHarmonize method, meaning that we do not propose a novel harmonization method but rather introduce an approach that internally integrates harmonization into ML pipelines while avoiding data leakage. This distinction is crucial, as the innovation lies not in altering the harmonization process but in structuring its application to ensure compatibility with ML workflows across both site-target dependence scenarios. Furthermore, this modular design of PrettYharmonize enables the use of other harmonization methods. Second, the output of PrettYharmonize is not a harmonized dataset but rather a final target prediction. Unlike traditional harmonization schemes, which produce a harmonized dataset for downstream use, our method uses harmonized data internally as an intermediate step to generate a prediction. By focusing on prediction, the method prioritizes practical utility in real-world applications where the primary goal is accurate target prediction rather than data transformation.
PrettYharmonize was rigorously validated on the MAREoS datasets, which were specifically designed to evaluate harmonization methods [10]. The strong performance observed on these benchmark datasets indicates that the internal harmonization scheme, based on pretended targets, successfully harmonizes data for the posterior use of the predictive model. In addition, on real-world datasets PrettYharmonize achieved competitive performance comparable to leakage-prone pipelines, demonstrating its practical utility. These results suggest that the proposed method is a promising alternative for real-world applications, particularly in settings where data leakage poses a substantial risk. Consequently, we advocate for its adoption in future use cases. Taken together, our findings underscore the need for meticulous evaluation when using data harmonization together with machine learning and encourage adoption of reproducible, open science practices to advance the field and benefit the wider community.
Conclusion
This study highlights three key findings regarding the use of harmonization methods in ML pipelines. First, while ComBat-based methods do not intrinsically cause data leakage, they might require that target variance be preserved as a covariate during training when a site-target relationship exists. As demonstrated, this requirement precludes their use in real-world ML applications, as it requires target values of the test samples that are not available. Our results empirically demonstrate that failing to preserve target variance leads to the removal of relevant signals, significantly undermining the effectiveness of harmonization. Second, harmonization did not lead to meaningful performance improvement when the site and targets are independent. This suggests that the benefits of harmonization may be limited in such cases, as the removal of site-related variability might not sufficiently enhance the signal-to-noise ratio for ML models. Finally, the proposed PrettYharmonize method advances the field by adapting neuroHarmonize for ML pipelines, effectively integrating harmonization while eliminating the need for test targets during training or transformation. This approach not only prevents data leakage by design but also achieved encouraging results, ensuring that the harmonization process remains both robust and practical for real-world applications.
@@ Conclusion
While this study provides valuable insights into harmonization methods, several limitations must be acknowledged. First, although our analysis focused on widely used ComBat-based techniques, other harmonization approaches, such as deep learning-based methods (e.g., style-matching generative models or variational autoencoders [1, 7]), were not explored. These methods may offer promising alternatives, particularly in complex scenarios where traditional approaches fall short, although they often require significant computational resources and large datasets.
Second, the impact of harmonization on feature selection and model interpretability has not been thoroughly investigated. Future research should explore how harmonization methods affect model behavior, especially in different domains and contexts, to better understand their broader implications for ML pipelines.
Third, while we simulated extreme site-target dependence and independence scenarios in a controlled manner, real-world cases are likely to fall somewhere along this spectrum. Our goal was to highlight potential problems that may arise from applying harmonization without proper consideration. Further studies are needed to investigate the effectiveness of harmonization methods at varying degrees of site-target dependence.
Finally, the proposed method introduces additional computational complexity compared to training a single neuroHarmonize model, as the “pretending” process requires multiple harmonization steps. However, the extra burden of this process is limited to applying a trained neuroHarmonize model several times (one harmonization transformation for each presented class), which is less time-consuming than training the model. Importantly, the computational cost of the pretending process does not scale with the number of sites to be harmonized, mitigating some of the practical limitations.
@@
All used MRI datasets are publicly available, possibly upon registration. The eICU dataset is publicly available at https://physionet.org/content/eicu-crd/2.0/ after registration. Registration includes the completion of a training course in research with human individuals at https://about.citiprogram.org/ and signing of a data use agreement mandating responsible handling of the data and adhering to the principle of collaborative research.
The PrettYharmonize library is publicly available at: https://github.com/juaml/PrettYharmonize. The scripts to replicate the experiments and to process the datasets are available at: https://github.com/juaml/harmonize_project
@@
To ensure the validity of PrettYharmonize, we benchmarked it in a classification problem using the datasets specifically designed to evaluate harmonization models [10]. This MAREoS dataset consists of eight datasets simulating 18 MRI features (cortical thickness, cortical surface area, or subcortical volumes). Four datasets contain a “True” signal and four only contain an Effect of Site (“EoS”) signal related to a binary target. In that sense, an ML model that learns the “EoS” signal can fraudulently achieve good classification performance. The signal, both the True or EoS, are called “Simple” and “Interactions”, depending on a linear or non-linear effect, respectively. Within each dataset, approximately 1000 samples, coming from 8 sites, were simulated. The datasets are provided as 10 train and test fold pairs. For the dataset containing only the EoS, the methods should be able to remove this effect and the classification performance should be at the chance level, i.e. balanced accuracy (bACC) of 50%. On the other hand, in the dataset with only the True signal, the harmonization models should not degrade the signal, and the bACC is expected to be a high value (bACC $\approx$ 80%).
@@
To empirically compare different harmonization schemes with and without site-target dependence, age regression, and sex classification were performed using MRI data. These targets were used as they are highly reliable and can be easily obtained. For all T1-weighted MR images, Voxel-Based Morphometry was performed using CAT12.8 [33] to obtain modulated gray matter (GM) volume, which was then linearly resampled to 8x8x8 mm3 voxels, resulting in 3747 voxels that were used as features. Five datasets were used: Amsterdam Open MRI Collection (AOMIC-ID1000) [34], The Enhanced Nathan Kline Institute (eNKI) [35], Cambridge Centre for Ageing Neuroscience (CamCAN) [36], 1000Brains [45], and the Southwest University Adult Lifespan Dataset (SALD) [37]. These datasets were selected as the data within each dataset was acquired only in one site thus avoiding additional confounding. The demographic information of these datasets is presented in Table 6.
@@
Forced site-target dependence.
Four datasets, AOMIC-ID1000, eNKI, CamCAN, and 1000Brains, were randomly subsampled in different age ranges, forcing a site-target dependence. The subsample was performed to ensure the same amount of subjects by each sex in each dataset (Table 7).
@@
Forced site-target independence.